V=4/3 pi r^3
Could I have the steps for solving for r^3? I need to get r^3 alone.

Answer by stanbon(75887)

V=4/3 pi r^3
Could I have the steps for solving for r^3? I need to get r^3 alone.
------------------
pi*r^3 = (3/4)V
--------
r^3 = (3/(4pi))V
---------------------------
Cheers,
Stan H.
--------------